At any point (x, y) on a curve, 𝑑^2𝑦/𝑑π‘₯^2 = 1-x^2


, and an equation of the tangent line to the


curve at the point (1, 1) is y = 2 – x. Find an equation of the curve.

1
Expert's answer
2022-02-08T07:28:01-0500

Solution:

"\dfrac{d^2y}{dx^2}=1-x^2
\\\Rightarrow \dfrac{dy}{dx}=x-\dfrac{x^3}3+C"

Given, tangent "y = 2 \u2013 x"

So, slope of tangent "=m=-1"

At (1, 1), "m=\dfrac{dy}{dx}"

"\Rightarrow x-\dfrac{x^3}3+C=m
\\\Rightarrow 1-\dfrac13+C=-1
\\ \Rightarrow C=-\dfrac53"

So, "\dfrac{dy}{dx}=x-\dfrac{x^3}3-\dfrac53"

Also, at any point, "m=\dfrac{dy}{dx}"

"x-\dfrac{x^3}3-\dfrac53=-1
\\ \Rightarrow x=-2,x=1"

For "x=-2: y=2-x=2+2=4"

For "x=1:y=2-x=2-1=1"

Points are (-2,4), (1,1).

Consider "\dfrac{dy}{dx}=x-\dfrac{x^3}3-\dfrac53" again.

On integrating:

"y=x^2-\dfrac{x^4}{12}-\dfrac53x+C_1"

Put (1,1) in this and solve for C1

"1=1-\dfrac{1}{12}-\dfrac53+C_1
\\\Rightarrow C_1=\dfrac74
\\ \therefore y=x^2-\dfrac{x^4}{12}-\dfrac53x+\dfrac74"